OpenStudy (anonymous):

Write e^4-2i in the a+bi form. Thank you

OpenStudy (anonymous):

\[e ^{(4-2i)}\] in the a+bi form.

geerky42 (geerky42):

HINT: \(\Large e^{xi} = \cos(x) + i\sin(x)\)

geerky42 (geerky42):

Better Hint: \(\Large e^{a+bi} = e^ae^{bi} = e^a(\cos(b) + i\sin(b))\)

geerky42 (geerky42):

@Haleo

geerky42 (geerky42):

Can you handle this problem now with my hints?

OpenStudy (anonymous):

@geerky42 , I will now try. I need a minute.

geerky42 (geerky42):

ok

OpenStudy (anonymous):

@geerky42, from my stuffing terms in the formula you gave (how is it called, I didn't go through it, I guess), I get the following: \[e ^{(4-2i)}= e ^{4}* e ^{-2i} = e ^{4}(\cos(-2) + i \sin (-2))\] But I don't feel it's right.

OpenStudy (anonymous):

Please, I need to finish calculating this! Help me, please.

OpenStudy (phi):

to put it into a+bi form, you should distribute the e^4 you can simplify cos(-2) to cos(2) and simplify sin(-2) to -sin(2)

OpenStudy (anonymous):

According to @ParthColi, a= \[e ^{4}\cos(-2) = e ^{4}\cos(2)\] and b = \[e ^{4}\sin(-2) = -e ^{4}\sin(2)\] But I don't understand how/why do the answers differ (both are -2) and how do get from that to the a+bi form.

OpenStudy (anonymous):

@phi

OpenStudy (phi):

you wrote **\[ e ^{(4-2i)}= e ^{4}* e ^{-2i} = e ^{4}(\cos(-2) + i \sin (-2)) \] which is the same as \[ e ^{4} \cos(-2) + i e ^{4}\sin (-2) \] it is well-known (to those who know trig well) that cos(-2)= cos(2) and sin(-2) is the same as -sin(2) so it looks a little nicer to write it as \[ e ^{4} \cos(2) - i\ e ^{4}\sin (2) \] notice the i in the 2nd term: you have the form a+b i you could write this as a decimal \[ \approx -22.72 -49.65 i \] but that is only an approximation

OpenStudy (anonymous):

pellet, it was correct when I tried to fondle a bit with the digits...the numbers seem so irrational that I thought it was a blah....Thank you for you help.

OpenStudy (anonymous):

hahah...they changed pellet to pellet...

OpenStudy (anonymous):

Oh, lol!